Prove that for no integer n, n6  +3n5_5n4_15n3+4n2+12n+3 is a perfect square.

Grade:10

1 Answers

Nishant Vora IIT Patna
9 years ago
Hello Student
Please fidn the soln

n6 +3n5_5n4_15n3+4n2+12n+3 =
= n5(n+3) –5n3(n+3)+4n(n+3) + 3
=(n+3)[n5 –5n3+4n] + 3
=n(n+3)[n4-5n2+4] + 3
=n(n+3)(n2-4)(n2-1) + 3
=(n-2)(n-1)n(n+1)(n+2)(n+3) + 3

So this can not be perfect square

Thansk